OpenStudy (stuck-help):

METAL AND FANSquare root x-2=2

OpenStudy (stuck-help):

@dan815 @perl

OpenStudy (anonymous):

x−2=2 step 1: Add 2 to both sides. x−2+2=2+2 x=4 Answer: x=4

OpenStudy (stuck-help):

X is squared